OpenStudy (anonymous):

A rectangle is to be drawn with perimeter 64 cm. If the length is to be 14 cm more than the width, determine the area of the rectangle. If I let x=length, y=width then would it be x=y+14 and y=x-1? or no?

OpenStudy (anonymous):

do you know what perimeter means?

OpenStudy (anonymous):

perimeter for a rectangle 2L+2W=P

OpenStudy (anonymous):

area of sqare (rectangle) A= lw

OpenStudy (anonymous):

oh wait okay i understand the question now

OpenStudy (anonymous):

so you just plug them in :P

OpenStudy (anonymous):

Then you distribute

OpenStudy (anonymous):

A rectangle is to be drawn with perimeter 64 cm. If the length is to be 14 cm more than the width, determine the area of the rectangle. If I let x=length, y=width then would it be x=y+14 and y=x-1? or no? ^ | is this your given?

OpenStudy (anonymous):

do you want me to plug them in or do you know how

OpenStudy (anonymous):

Plug L=w+14 into 2L+2W=64

OpenStudy (anonymous):

2L+2W=P 2(14)+2y=64 28+2y=64 -28 -28 2y=36 / / 2 2 y=18 A=14 * 18 A=252?? I know that the answer is wrong cause the back of the book says 207 cm^2. What step did I make a mistake?

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

perimeter means the total measurement of all sides of a figure